# A Plug-and-Play Method for Guided Multi-contrast MRI Reconstruction based on Content/Style Modeling

### v1

@@ IV-B Benchmark and Ablation
TABLE II: Data split for benchmarking. Splits were performed at subject-level, each subject belonging to exactly one subset.
@@
NYU model-train 200 200 200 / 200
model-val 27 27 27 / 27
recon-val 50 50 50 / 50
recon-test 50 50 50 / 50
Total 327 327 327 / 327
LUMC-TRA model-train 295 418 360 / 415
model-val 16 17 17 / 17
@@
LUMC-COR model-train 242 277 269 / 272
model-val 18 18 18 / 18
@@
The NYU brain dataset includes raw data and DICOM scans of 4 contrasts – T1W with and without gadolinium agent, T2W, and FLAIR. Following [33], a DICOM subset of 327 subjects was obtained based on T2W and non-gadolinium T1W scans. Single-coil T2W k-space data was simulated via Fourier transform with 1D Cartesian random sampling of $R\in\{2,3,4,5\}$ and added Gaussian noise of $\sigma=0.01\mathrm{max}(x_{2}^{*})$.
Our in-house datasets consisted of brain scans of patients from LUMC, the use of which was approved for research purpose by the institutional review board. A total of 1669 brain scans were obtained from 817 clinical MR examinations of 630 patients acquired on 3T Philips Ingenia scanners. We focused on accelerating two T2W sequences – (a) 2D T2W TSE transversal (FA=90°, TR=4000-5000ms, TE=80-100ms, ETL=14-18, median voxel size=0.37$\times$0.5mm2, median FOV=238$\times$190mm2, median slice thickness=3mm, median slices=50), and (b) 2D T2W TSE coronal (FA=90°, TR=2000-3500ms, TE=90-100ms, ETL=17-19, median voxel size=0.39$\times$0.47mm2, median FOV=130$\times$197mm2, median slice thickness=2mm, median slices=15). For guidance, two corresponding T1W sequences were used – (a) 3D T1W TFE transversal (FA=8°, TR=9.8-9.9ms, TE=4.6ms, ETL=200, median voxel size=0.98$\times$0.99$\times$0.91mm3, median FOV=238$\times$191$\times$218mm3), and (b) 3D T1W TSE coronal (FA=80-90°, TR=500-800ms, TE=6.5-16ms, ETL=10-13, median voxel size=0.59$\times$0.62$\times$1.19mm3, median FOV=130$\times$238$\times$38mm3). The transversal and coronal protocols were considered as two separate datasets, namely LUMC-TRA and LUMC-COR. The multi-coil T2W raw data comprised of 6-channel (LUMC-TRA) and 13-channel (LUMC-COR) k-space and coil sensitivity maps. This k-space was already undersampled (1D Cartesian random) at acquisition with clinical acceleration of $R$=1.8-2. We further undersampled it retrospectively to higher accelerations $R\in\{4,6,8,10\}$ by dropping subsets of the acquired lines.
Each dataset was split into 4 subsets – (a) model-train, (b) model-val, (c) recon-val, and (d) recon-test. The former two were used to train and tune the plug-and-play models, recon-val to tune the reconstruction algorithms, and recon-test as the held-out test set. An overview of the data split is shown in Table II. Additionally, 20 subjects from model-train were designated for PFT. Spatially aligned images required by all except the unpaired training set were obtained via rigid registration using Elastix [38]. We empirically found the ContentSize-1 and ContentSize-2 configurations as optimal for the NYU and LUMC datasets, respectively, which reflects the lower effective content and hence greater difficulty of the LUMC datasets. For each dataset, the model was pre-trained and fine-tuned for 400k and 50k iterations, respectively. We used 3 perceptual metrics for evaluation: SSIM, HaarPSI, and DISTS. While SSIM is commonly used, HaarPSI and DISTS are known to correlate better with visual judgment of image quality [39]. All three metrics are bounded to $[0,1]$ with $1$ representing perfect image quality. We additionally conducted paired Wilcoxon signed-rank tests to measure statistical significance when comparing pairs of algorithms.
Fig. 5 plots the benchmark metrics, which can be summarized in the following three trends. First, pure image translation was worse compared to single-contrast reconstruction, especially at lower acceleration factors (comparing with PnP-CNN, $p$$<$0.05 for all metrics and datasets). Combining it with L2-regularized reconstruction improved SSIM and HaarPSI ($p$$<$0.05 for all datasets and accelerations except SSIM at $R$=3 in NYU), but not necessarily DISTS, suggesting that the the available complementary information was not fully utilized. PnP-MUNIT was consistently better than both image translation and naive guidance ($p$$<$0.05 throughout for both cases). It also outperformed the conventional guided CS-STV ($p$$<$0.05 throughout, except SSIM at $R$=4 in LUMC-COR). Second, both PnP-CNN and PnP-MUNIT performed similarly at lower acceleration. At higher acceleration, PnP-MUNIT outperformed PnP-CNN ($p$$<$0.05 for all metrics and datasets except SSIM at $R$=8 in LUMC-COR). Third, introducing PFT and CR into the ablated PnP-MUNIT improved the reconstructions ($p$$<$0.05 throughout in both cases except in DISTS for the latter at $R$=10 in LUMC-TRA). The contribution of PFT was almost constant across $R$, whereas that of CR decreased with $R$, which was expected since CR depends on the measured k-space data to refine the content.
@@
Figure 6: Sample slice from the LUMC-COR recon-test set comparing PnP-MUNIT with PnP-CNN and pure image translation.
@@
Fig. 6 shows an example from LUMC-COR representative of the first two trends. Image translation produced severe anatomical defects and predicted structure in the lower region of the image where the ground truth contains low signal. At $R$=4, PnP-MUNIT was visually similar to PnP-CNN, with the main difference being a blur effect in PnP-CNN and mild texture artifacts in PnP-MUNIT. At $R$=10, PnP-CNN reconstruction was non-viable, containing severe blur and artifacts (green arrows). On the other hand, PnP-MUNIT did not degrade much from the $R$=4 case and most fine structures were sharply resolved (green arrows). However, an interesting failure mode of PnP-MUNIT at high acceleration was the subtle localized distortions in the anatomy (pink arrows).
IV-C Radiological Evaluation
@@
We conducted the radiological evaluation on a small sample of LUMC-TRA recon-test set. We selected 3 cases that showed brain metastases in corresponding gadolinium-enhanced T1W scans. The evaluation comprised two parts, namely visual quality and pathology. Four visual quality criteria were used, namely sharpness, noise, artifacts, and contrast between gray and white matter and CSF. For pathology, we used three criteria, namely the number and sharpness of hyperintense areas within or surrounding metastases and the overall diagnostic quality of the scan for brain metastases. The images were scored using a five-point Likert scale [40] – (1) non-diagnostic, (2) poor, (3) fair, (4) good, and (5) excellent diagnostic quality. We considered PnP-MUNIT and PnP-CNN at 4 clinically realistic accelerations $R\in\{3,4,5,6\}$. In total, 9 images per patient, including the reconstructions and the clinical ground truth of $R$=2, were presented to a senior neuroradiologist who scored each image individually, blinded to the reconstruction method and $R$. Figure 7 plots the evaluation result. At $R$=3, both algorithms produced at least “fairly diagnostic” reconstructions and allowed the detection of more than 90% of metastases. PnP-CNN was better than PnP-MUNIT at $R\in\{3,4\}$ in terms of sharpness, contrast, and the pathology criteria. At $R$=5, PnP-MUNIT matched or exceeded PnP-CNN in terms of visual quality, although being worse in terms of pathology criteria. At $R$=6, PnP-CNN sharply dropped to non-diagnostic level, whereas PnP-MUNIT was better, especially in terms of sharpness and contrast as well as in overall diagnostic quality.
V Discussion
We explored the interpretability aspect of PnP-MUNIT in terms of the effective content in a two-contrast dataset (Section IV-A1) and obtained a notion of limit to guided reconstruction in terms of the reference image resolution. This limit would indeed depend on more fundamental factors such as MR sequence types, which we did not investigate here. Our data-driven definition of content and style is loosely analogous to MR physics-based descriptions of quantitative tissue maps and acquisition-related factors. Hence, a further level of interpretability could be achieved by constraining the learned content to represent physically meaningful tissue parameters. Content discrepancy is characterized as a limiting factor in guided reconstruction (Section III-C) and plays a major role throughout this work. Viewed as a set of quasi tissue maps, estimating content from a single image is challenging, ideally requiring a time series of images, which captures the magnetization dynamics. Content discrepancy is, therefore, inevitable. We proposed two complementary ways to minimize its effect on the reconstruction. While paired fine-tuning improves the model using a small amount of paired data, content refinement (CR) corrects for the remaining discrepancy adaptively during reconstruction using the measured k-space. A drawback of CR, however, is slower convergence (Section IV-A2). This could be mitigated by partially adopting an unrolled network design. The resulting tradeoff between model latency and generalizability would be interesting to explore.

### v2

@@ IV-A Experiment Design
TABLE II: LUMC data split. The splits were made at the subject-level, i.e. each subject belonged to exactly one subset. Splits containing only image-domain data are marked with I.
@@
LUMC-TRA model-trainI 295 418 360 / 415
model-valI 16 17 17 / 17
@@
LUMC-COR model-trainI 242 277 269 / 272
model-valI 18 18 18 / 18
@@
IV-A1 Simulation
The goal of our simulation experiments was to empirically test several properties of our algorithm under highly controlled settings. To this end, we used simulated T1W and T2W images based on BrainWeb phantoms [44]. BrainWeb provides 20 anatomical models of normal brain, each comprised of fuzzy segmentation maps of 12 tissue types. The 20 volumes were first split in 18:1:1 ratio for model training, validation, and reconstruction testing, respectively. T1W/T2W spin-echo scans were simulated using TE/TR values randomly sampled from realistic ranges. For testing the reconstruction, 2D single-coil T2W k-space data was simulated via Fourier transform and 1D Cartesian random sampling at various accelerations. We analyzed the effect of the content/style model’s disentanglement level and its content capacity on the reconstruction quality, the convergence of the PnP-MUNIT algorithm, and the effectiveness, robustness, and tuning of the CR module.
IV-A2 Benchmark on NYU DICOM dataset
In the NYU benchmark, we compared PnP-MUNIT against end-to-end reconstruction methods. The question we sought to answer was whether or not PnP-MUNIT, which requires only image-domain training data, can outperform methods that additionally require k-space training data. The NYU brain dataset [45, 46] includes raw data and DICOM scans of 4 contrasts – T1W with and without gadolinium agent, T2W, and FLAIR. Following [41], a paired DICOM subset of 327 subjects was obtained based on T2W and non-gadolinium T1W scans. All T1W scans were rigid-registered with the corresponding T2W scans. Single-coil T2W k-space data was simulated via Fourier transform with 1D Cartesian random sampling at $R\in\{2,3,4,5\}$ and added Gaussian noise of $\sigma=0.01\mathrm{max}(x_{2}^{*})$. The dataset was split into 4 subsets – model-train (200), model-val (27), recon-val (50), and recon-test (50). Images from the former two splits were used to pre-train, fine-tune, and validate the MUNIT model. On the other hand, end-to-end models were trained and validated on the aligned T1W, the simulated T2W k-space, and the T2W ground-truth from these two splits. The recon-val split was used to tune the PnP-MUNIT algorithm, and recon-test was the held-out test set. We benchmarked PnP-MUNIT against one well-known single-contrast network – MoDL [4] – and two recent multi-contrast networks – MTrans [32] and MC-VarNet [15].
IV-A3 Benchmark and ablation on LUMC multi-coil dataset
In our benchmark on clinical multi-coil data, our goal was to test PnP-MUNIT on a constrained real-world problem where only the image-domain data was available for training. This is often the case, as in clinical practice raw data is discarded after acquisition and only the final reconstructions are retained. Moreover, the T1W/T2W images were not fully subject-wise paired, representing a realistic case of data imbalance. Our in-house data consisted of brain scans of patients from LUMC, the use of which was approved for research purpose by the institutional review board. A total of 1669 brain scans were obtained from 817 clinical MR examinations of 630 patients acquired on 3T Philips Ingenia scanners. We focused on accelerating two T2W sequences – (a) 2D T2W TSE transversal and (b) 2D T2W TSE coronal. For guidance, two corresponding T1W sequences were used – (a) 3D T1W TFE transversal and (b) 3D T1W TSE coronal. The transversal and coronal protocols were considered as two separate datasets, namely LUMC-TRA and LUMC-COR. Table I shows an overview of the four sequences. Note that the in-plane resolution of the T1W scans was 1.3-2.5 times as low as that of the T2W scans, making these datasets more challenging for T1W-guided T2W reconstruction.
As with the NYU dataset, the LUMC datasets were split into model-train, model-val, recon-val, and recon-test. The former two splits contained only image-domain data, which was used to train and validate MUNIT. The MUNIT models were pre-trained on the full model-train splits, ignoring any pairing between T1W and T2W scans. Additionally, 20 subjects from the model-train split were designated for PFT where the pairing information was used in training and the reference scans were aligned via registration. The recon-val and recon-test splits additionally included multi-coil T2W raw data, which comprised 6-channel (LUMC-TRA) and 13-channel (LUMC-COR) k-space and coil sensitivity maps. This k-space was already undersampled (1D Cartesian random) at acquisition with clinical acceleration of $R$=1.8-2. We further undersampled it retrospectively to higher accelerations $R\in\{4,6,8,10\}$ by dropping subsets of the acquired lines. An overview of the data split is shown in Table II. Spatially aligned images required by all but the unpaired training set were obtained via rigid registration and the reference T1W images were resampled to the T2W resolution. During unpaired training, all images were resampled to the median T2W resolution.
Given the of absence of k-space training data, end-to-end reconstruction methods were not feasible as baselines. Hence, we compared PnP-MUNIT with only the feasible types of baselines, i.e. classical, semi-classical, and plug-and-play reconstruction and image-to-image translation. Among classical methods, we used the unguided L1-wavelet CS (CS-WT) and the guided STV-based CS (CS-STV) [7]. As an unguided plug-and-play baseline, we used PnP-CNN [23], and as a representative image translation baseline, we compared against MUNIT itself. Using MUNIT, deterministic image translation was approximated by combining the reference content with 200 randomly sampled T2W style codes and taking a pixel-wise mean of the resulting synthetic images to obtain a single synthetic image. As semi-classical guided baseline, we used (PROSIT) [39], which combines deterministic image translation with L2-regularized least-squares reconstruction. Additionally, as an ablation study for PnP-MUNIT, we ablated PFT and CR to assess their contribution, and finally, as an upper-bound for PnP-MUNIT representing zero content discrepancy, we disabled PFT and CR and used the ideal content $c^{*}$. In both NYU and LUMC benchmarks, we used 3 perceptual metrics for evaluation – SSIM, HaarPSI, and DISTS. While SSIM is used commonly, HaarPSI and DISTS are known to correlate better with visual judgment of image quality [47]. All three metrics are bounded in $[0,1]$ where $1$ represents perfect image quality. We conducted paired Wilcoxon signed-rank tests to measure statistical significance when comparing pairs of algorithms.
IV-A4 Radiological evaluation
Finally, as an extension to the LUMC benchmark, we conducted a radiological evaluation assessing the visual and diagnostic quality of PnP-MUNIT reconstructions. The study was conducted on a small sample of the LUMC-TRA recon-test set. With the help of a junior radiologist, we selected 3 cases with brain metastases. The evaluation comprised two parts, namely visual quality and pathology. Four visual quality criteria were used, namely sharpness, noise, artifacts, and contrast between gray and white matter and CSF. For pathology, we used three criteria, namely the number and sharpness of hyperintense areas within or surrounding metastases and the overall diagnostic quality of the scan for brain metastases. The images were scored using a five-point Likert scale [48] – (1) non-diagnostic, (2) poor, (3) fair, (4) good, and (5) excellent diagnostic quality. A diagnostic-quality reconstruction was defined as one that showed at least 90% of the metastases and scored at least “fairly diagnostic quality” on all other criteria. [^2] In clinical practice, while the remaining 10% of the metastases may not be detected in the reconstructed T2W scan, they would be prominently visible in the corresponding gadolinium-enhanced T1W scan, and hence would not go undetected. We evaluated PnP-MUNIT and PnP-CNN at 4 clinically realistic accelerations of $R\in\{3,4,5,6\}$. In total, 9 images per patient, including the (4$\times$2=8) reconstructions and the clinical ground truth (acquired at $R$=2 and reconstructed by the vendor software), were presented to a senior neuroradiologist who scored each image individually, blinded to the reconstruction method and $R$.
IV-B Implementation Details
We used the same general residual architecture for MUNIT encoders and decoders as the original paper, except with an additional layer at the end of the content encoder to produce content maps of given number of channels. We used 2 or 4 channels, depending on the content downsampling factor. The discriminators were implemented as multi-scale PatchGAN networks enabling them to locally assess the input images for realism at different scales. We used 1 scale in the BrainWeb simulations and 3 scales in the NYU and LUMC benchmark. We additionally conditioned the discriminators on foreground masks of the images to penalize background signal. To stabilize GAN training, we used spectral normalization [49]. For hyperparameters, we found that pre-training loss weight values $\alpha_{1}=\alpha_{2}=\alpha_{3}=1$ and paired fine-tuning values $\beta_{1}=\beta_{2}=\beta_{3}=1$ worked generally well for the in vivo datasets. In the benchmarks, the CR module parameter $\gamma$ of PnP-MUNIT was tuned per acceleration for each dataset and the number of iterations was set to 200.
In the simulations, all MUNIT models were pre-trained for 200k and were not fine-tuned for the sake of simplicity, unless stated otherwise in Section V-A. In both NYU and LUMC benchmarks, the MUNIT models were pre-trained for 400k iterations and fine-tuned for 50k iterations. All software was implemented in PyTorch, and image registration was performed using Elastix [50]. All training runs were performed on a compute node with an NVIDIA Quadro RTX 6000 GPU.
V Results
TABLE III: Reconstruction PSNR (dB) over the BrainWeb test volume (300 slices). PSNR worse than L1-wavelet CS is marked as .
Disentanglement Strength R=2 R=4
$\alpha_{2}=\alpha_{3}=10$ 15.14 $\pm$ 0.06 12.84 $\pm$ 0.05
$\alpha_{2}=\alpha_{3}=1$ 28.43 $\pm$ 0.07 26.13 $\pm$ 0.07
$\alpha_{2}=\alpha_{3}=0.1$ 29.84 $\pm$ 0.05 27.33 $\pm$ 0.05
$\alpha_{2}=\alpha_{3}=0.01$ 25.34 $\pm$ 0.06 23.03 $\pm$ 0.05
$\alpha_{2}=\alpha_{3}=0.001$ 18.17 $\pm$ 0.07 15.88 $\pm$ 0.05
TABLE IV: Reconstruction PSNR (dB) over the BrainWeb test volume (300 slices) at $R$=4. Optimal content capacity is shown in bold. PSNR worse than L1-wavelet CS is marked as .
Model Config Data Config
RefRes-1 RefRes-2 RefRes-4 RefRes-8
ContentRes-1 27.90$\pm$0.08 21.17$\pm$0.05 19.11$\pm$0.07 16.43$\pm$0.07
ContentRes-2 26.19$\pm$0.08 22.80$\pm$0.08 20.05$\pm$0.08 16.57$\pm$0.08
ContentRes-4 23.98$\pm$0.08 20.88$\pm$0.10 19.44$\pm$0.09 16.88$\pm$0.06
Figure 4: Effectiveness of the CR module in resolving contrast-specific structure, which, in this case, was a simulated lesion.
V-A Experiments on Simulated MR Datasets
V-A1 Perturbation of disentanglement strength
PnP-MUNIT relies heavily on a sufficient disentanglement of content and style from the images. Here, we evaluated the effect of perturbing the disentanglement loss weights $\alpha_{2}$ and $\alpha_{3}$ (Eq. (6)) of MUNIT on the PnP-MUNIT reconstruction quality. Table III shows the reconstruction PSNR for different weight values. We observed that optimal disentanglement, and thus the best reconstruction, was achieved at $\alpha_{2}=\alpha_{3}=0.1$. At lower values, poor disentanglement of content and style in the model resulted in poor PnP-MUNIT reconstruction quality. This was expected since PnP-MUNIT relies on the assumption that the content representation is (sufficiently) contrast-independent, which was less enforced at these levels. At higher values, the training over-emphasized disentanglement while under-emphasizing the GAN and image recovery loss terms, thereby leading to worse preservation of image information in the latent space and thus, to worse reconstructions.
V-A2 Content resolution analysis
PnP-MUNIT relies on a reference image of sufficiently high resolution to provide guidance to the reconstruction. The goal of this experiment was to determine the effect of lowering the reference image resolution on the reconstruction quality and using the content/style model’s content capacity (i.e. the spatial resolution of the content relative to the images) to explain this effect. We simulated 4 datasets, denoted as RefRes-$n$, where $n\in\{1,2,4,8\}$ represents the reference domain downsampling factor. In the $n$=1 case, T1W/T2W images had the same resolution (as that of the underlying tissue maps), whereas in the subsequent cases, T1W images were blurred to contain only the lower $1/n$ frequency components, while maintaining the same spatial resolution. For each dataset, we trained 3 content/style models with content capacity denoted as ContentRes-$m$, where $m\in\{1,2,4\}$ is the content downsampling factor, which depends on the number of downsampling blocks in the networks. E.g. ContentRes-2 model produced content maps half the spatial resolution of the ContentRes-1 model.
Figure 5: The effect of misaligned reference on the reconstruction and the robustness of the CR module to it.
Figure 7: Evaluation plots over recon-test subset of NYU DICOM dataset consisting of 550 slices.
Figure 8: Two examples from the NYU benchmark. (a) Comparison of all the methods at the highest acceleration of $R$=5. (b) Comparison between PnP-MUNIT and MC-VarNet at $R$=2 and $R$=4 on a pathological case where many features of the edema are T2W contrast-specific.
Table IV compares PnP-MUNIT reconstruction quality across these configurations. PFT and CR were disabled here for simplicity. We observe two trends. First, the reconstruction quality generally decreases from left to right, eventually dropping below L1-wavelet CS reconstruction, suggesting a decrease in the amount of shared local information contained in the reference contrast. Second, the optimal content capacity of the model decreases with the reference resolution in accordance with the actual amount of this shared information. In other words, this optimal content capacity corresponds to the amount of contrast-independent structure the model discovers in the dataset, e.g. RefRes-1 dataset had 4 times as much shared structure (due to twice as large optimal content resolution) as RefRes-2 and RefRes-4 datasets and hence, the RefRes-1 dataset was significantly more effective in guided reconstruction. On the other hand, Refsize-2 and RefRes-4 datasets had similar levels of shared structure despite the lower reference resolution in the latter, suggesting a more complex relationship between the reference resolution and the shared content. In the RefRes-8 case, PnP-MUNIT dropped to a similar level of quality as conventional CS, suggesting a lowerbound on the amount of shared information for PnP-MUNIT to be effective, specifically 1/42=1/16 times the local information of the full-resolution case.
V-A3 Convergence
We explored the convergence of PnP-MUNIT by comparing it with two variants. The first used true content $c^{*}$ of the ground truth T2W image, thus assuming zero content discrepancy and representing an upper-bound of PnP-MUNIT. The second used reference content $\hat{c}$ with CR step disabled, representing a lower-bound of PnP-MUNIT where the non-zero content discrepancy is not corrected. Here and in the following simulation experiments, we used the RefRes-1 dataset and the ContentRes-1 model additionally fine-tuned on 2 training volumes for 50k iterations. Fig. 3 shows the convergence curves and intermediate reconstructions. The upper-bound version, given $c^{*}$, converged in a single iteration, while with $\hat{c}$, the ablated version converged equally fast but to a sub-optimal solution. Enabling the CR step closed the gap with the upper-bound, although costing convergence rate.
V-A4 Effectiveness in resolving contrast-specific structure
In order to test the effectiveness of the CR module in resolving structure present in the target contrast but absent in the reference contrast, we simulated a lesion in the T2W image and performed reconstruction with and without the CR module. As shown in Fig. 4, the CR module was able to recover the lesion fully at $R$=2 and substantially at $R$=4.
V-A5 Sensitivity to misalignment of the reference
While PnP-MUNIT expects an aligned reference image at reconstruction time, the CR module should, in principle, correct for small misalignments. To test this, we simulated a 2 rotation in the reference image and performed reconstruction with and without the CR module, comparing also with PnP-MUNIT that used an aligned reference image. As shown in Fig. 5, although the reconstruction was sensitive to the misalignment in the absence of CR, it significantly improved with CR enabled.
Figure 9: Evaluation plots over recon-test subset of LUMC-TRA and LUMC-COR datasets consisting of 1366 and 200 slices, respectively.
@@ V-A5 Sensitivity to misalignment of the reference
Figure 10: Sample slice from the LUMC-COR recon-test set comparing PnP-MUNIT with PnP-CNN and MUNIT image translation.
V-A6 Tuning the CR step size
The hyperparameter $\gamma$ in Algorithm 1 controls the CR strength. Fig. 6 shows reconstruction quality as a function of acceleration and k-space noise level, where we observe that the optimal $\gamma$ decreased with the amount and quality of k-space data. This makes sense as enforcing agreement with the k-space gradually becomes less beneficial, thereby contributing less to reconstruction quality.
V-B Benchmark on NYU DICOM Data
Fig. 7 shows evaluation metrics for PnP-MUNIT and the end-to-end baselines on the NYU DICOM dataset. Despite being trained only on image-domain data, PnP-MUNIT largely outperformed MoDL ($p<0.05$ throughout, except SSIM at $R$=2 and DISTS at $R\in${2,3,4}) and MTrans ($p<$0.05), and was roughly comparable to MC-VarNet. Moreover, as shown in Fig. 8a, PnP-MUNIT could resolve certain fine details even at the high acceleration of $R$=5, which the baselines failed at. In the pathological case shown in Fig. 8b, PnP-MUNIT produced sharper reconstructions preserving the structure of the edema better than MC-VarNet despite lower metrics when many features of edema were absent in the reference image.
V-C Benchmark and Ablation on LUMC Multi-coil Data
The LUMC datasets represent a more challenging problem with real-world data constraints. Additionally, in the light of the content resolution analysis (Section V-A2), we empirically found the ContentRes-2 model configuration as optimal for the LUMC datasets, compared to NYU DICOM data for which ContentRes-1 was the optimal model configuration. This reflects the lower effective content in LUMC image data and hence a greater difficulty for guided reconstruction.
Fig. 9 plots the benchmark metrics for the LUMC test sets, which can be summarized in the following three trends. First, pure image translation with MUNIT was worse compared to single-contrast reconstruction, especially at lower acceleration factors (comparing with PnP-CNN, $p$$<$0.05 for all metrics and both datasets). Combining it with L2-regularized reconstruction in PROSIT improved SSIM and HaarPSI ($p$$<$0.05 for both datasets and accelerations), but not necessarily DISTS, suggesting that the the available complementary information was not fully utilized. PnP-MUNIT was consistently better than both MUNIT image translation and PROSIT ($p$$<$0.05 throughout for both cases). It also outperformed the conventional guided CS-STV ($p$$<$0.05 throughout, except SSIM at $R$=4 in LUMC-COR). Hence, using both measured k-space and reference scan via PnP-MUNIT was more beneficial than (a) using either one of them and (b) combining both using hand-crafted priors, suggesting that our approach maximally exploits the complementary information. Second, both PnP-CNN and PnP-MUNIT performed similarly at lower acceleration, except in terms of SSIM and HaarPSI at $R$=4 and $R$=6 on LUMC-COR where PnP-CNN was slightly better. At higher acceleration, PnP-MUNIT outperformed PnP-CNN ($p$$<$0.05 for all metrics and both datasets except SSIM at $R$=8 in LUMC-COR). Compared to PnP-CNN, PnP-MUNIT allowed up to 32.6% more acceleration for a given SSIM. [^3] Based on linear interpolation of LUMC-TRA SSIM, PnP-MUNIT and PnP-CNN allowed $R$=10 and $R$=6.7, the maximum difference in $R$, at median SSIM=0.906. Hence, a difference of 32.6%. Third, introducing PFT and CR into the ablated PnP-MUNIT improved the reconstructions ($p$$<$0.05 throughout in both cases except in DISTS for the latter at $R$=10 in LUMC-TRA). The contribution of PFT was almost constant across $R$, whereas that of CR decreased with $R$, which was expected since CR depends on the measured k-space data to refine the content.
@@ V-C Benchmark and Ablation on LUMC Multi-coil Data
Fig. 10 shows an LUMC-COR example representative of the first two trends. MUNIT image translation produced severe anatomical defects and predicted false structure in the lower region of the image where the ground truth, in fact, contains low signal. At $R$=4, PnP-MUNIT was visually similar to PnP-CNN, with the main difference being a blur effect in PnP-CNN and mild texture artifacts in PnP-MUNIT. At $R$=10, PnP-CNN reconstruction was non-viable, containing severe blur and artifacts (green arrows). On the other hand, PnP-MUNIT did not degrade much from the $R$=4 case and most fine structures were sharply resolved (green arrows). However, an interesting failure mode of PnP-MUNIT at high acceleration was the subtle localized distortions in the anatomy (pink arrows).
@@
V-D Radiological Evaluation
Figure 11 plots the radiological evaluation result. At $R$=3, both algorithms produced at least “fairly diagnostic” reconstructions and allowed the detection of more than 90% of metastases, a difference of 33.3% of k-space samples over the clinical reconstructions of $R$=2. PnP-CNN was better than PnP-MUNIT at $R\in\{3,4\}$ in terms of sharpness, contrast, and the pathology criteria. At $R$=5, PnP-MUNIT matched or exceeded PnP-CNN in terms of visual quality, although being worse in terms of pathology criteria. At $R$=6, PnP-CNN sharply dropped to non-diagnostic level, whereas PnP-MUNIT was better, especially in terms of sharpness and contrast as well as in overall diagnostic quality.
VI Discussion
In this work, we modeled the two-contrast MR image data in terms of two latent generative factors – content representing the contrast-independent structure and style representing the contrast-specific variations. Our data-driven definitions of content and style, though seemingly related to the MR physics-based representations of quantitative maps and acquisition-related factors, are rather nebulous compared to precisely defined physical concepts. For instance, MR quantitative maps are theoretically objective representations of tissues in terms of physical variables such as relaxation times and are ideally independent of the contrast-generating sequence parameters. On the other hand, our content and style representations are only defined statistically in terms of the given two contrasts and are, thus, neither as objective nor as independent in the ideal physical sense. That being said, we would argue that there is potential value in augmenting the data-driven content/style model with MR physical models, e.g. by constraining the learned content to represent physically meaningful anatomical properties, introducing elements of a physical model (e.g. Bloch equation) into the decoder network, etc., to enhance interpretability and reconstruction quality.
Optimal content capacity was defined as a quantity representing the amount of shared local information contained in a two-contrast dataset. In terms of this quantity, we analyzed the effect of reference image resolution on reconstruction quality, obtaining a lowerbound for PnP-MUNIT to be effective (Section V-A2). This quantity would depend on more fundamental factors such as MR sequence types, which we did not investigate here. While we limited our experiments to T1W and T2W sequences, PnP-MUNIT is applicable to any pair of contrasts, with the image quality limited mainly by the amount of structural information shared between them. Future work could help empirically determine which sequence pairs are more amenable to guided reconstruction than others. Note that for a different pair of contrasts, say PD-weighted and T2W, a fresh content/style model needs to be trained since content and style are defined in the context of the specific contrast pair. It is, in principle, also possible to model the content and style for more than two contrasts. Given number of contrasts $N$, content is defined as the local structure underlying all $N$ contrasts, while style encodes intra-contrast variations of each individual contrast. This model would include $N$ sets of encoders and decoders and would require an (unpaired) image-domain training dataset of these $N$ contrasts. Subsequently, our PnP-MUNIT framework can be extended to use multiple reference contrasts to guide a given target contrast. The specific question to investigate then would be – given $P$ reference contrasts (where $1<P<N$), how to aggregate their contents such that the prior structural information about the target image is maximized?
Content discrepancy was characterized as a quantity representing the gap between the true content of the target image and our estimation of this content, which is another limiting factor in PnP-MUNIT reconstruction. In MUNIT, the decoders model the (forward) generative process that maps the latent variables (i.e. content and style) to the observable variable (i.e. images). Ideally, the encoders must be a perfect inverse of this forward process. However, estimating the multi-channel contrast-independent content from a single image (using the content encoders) is a challenging (and perhaps ill-conditioned) inverse problem. Hence, some errors are to be expected in the content estimated from the reference image. Moreover, the reference image may contain unfavorable differences from the target image, such as misalignment or artifacts, which introduce more errors. Therefore, content discrepancy is expected to be a non-zero quantity. We proposed two complementary ways to minimize its effect on the reconstruction. While paired fine-tuning improves the model using a small amount of paired image-domain training data, content refinement (CR) aims to correct the remaining discrepancy during reconstruction using the measured k-space. A drawback of the CR module, however, is slower convergence (Section V-A3). For example, the run-time for PnP-MUNIT on the GPU was around 17 seconds for a slice of matrix size 349$\times$284, as compared to 6 seconds for CS-WT, 1.4 seconds for PnP-CNN, and less than 0.1 second for MC-VarNet. This latency issue could be mitigated by approximating the iterative process with an unrolled network design, although perhaps at a cost of the generalizability offered by the plug-and-play design. The resulting trade-off between model latency and generalizability would be interesting to explore in the future.